# Partial-Label Learning with a Reject Option

### v3

@@ Abstract
In real-world applications, one often encounters ambiguously labeled data, where different annotators assign conflicting class labels. Partial-label learning allows training classifiers in this weakly supervised setting, where state-of-the-art methods already show good predictive performance. However, even the best algorithms give incorrect predictions, which can have severe consequences when they impact actions or decisions. We propose a novel risk-consistent partial-label learning algorithm with a reject option, that is, the algorithm can reject unsure predictions. Extensive experiments on artificial and real-world datasets show that our method provides the best trade-off between the number and accuracy of non-rejected predictions when compared to our competitors, which use confidence thresholds for rejecting unsure predictions instead. When evaluated without the reject option, our nearest neighbor-based approach also achieves competitive prediction performance.
@@ 1 Introduction
However, even the best algorithms give incorrect predictions. These errors can have severe consequences when they impact actions or decisions. Consider, for example, safety-critical domains such as the classification of medical images (Yang et al., 2009; Lambrou et al., 2011; Senge et al., 2014; Kendall & Gal, 2017; Reamaroon et al., 2019) or the control of self-driving cars (Xu et al., 2014; Varshney & Alemzadeh, 2017; Hubmann et al., 2017; Shafaei et al., 2018; Michelmore et al., 2020). One option to limit fallacies is to refuse to give an individual prediction if unsure and, instead, let humans decide on the label of an instance or the actions to take. A reject option allows one to abstain from certain predictions based on a threshold. Naturally, there is a trade-off arising between the number and accuracy of non-rejected predictions. In the supervised setting, reject options have already been studied, both, for classification (Hellman, 1970; Herbei & Wegkamp, 2006) and regression tasks (Zaoui et al., 2020). In PLL, a reject option can also mitigate misclassifications, which are more likely due to the weak supervision. We are the first to study reject options in PLL.
Determining whether to reject a prediction of a PLL algorithm is highly non-trivial as the uncertainty involved in prediction-making originates from multiple sources (Hora, 1996). On the one hand, (i) there is inherent uncertainty due to ambiguously labeled data (aleatoric uncertainty). This ambiguity is influenced by both instances and classes: In some regions of the instance space, annotators are more likely to mislabel instances than in others. Additionally, some class labels are inherently more similar than others. On the other hand, (ii) there is uncertainty due to the lack of knowledge regarding the relevance of each candidate label (epistemic uncertainty). Optimally, an algorithm does not mix (i) and (ii) in deciding whether to reject a prediction.
In fact, Fisher (1930, 1935) already notes that “not knowing the chance of mutually exclusive events [epistemic uncertainty] and knowing the chance to be equal [aleatoric uncertainty] are two quite different states of knowledge.” One can reduce epistemic uncertainty with more data or better models; aleatoric uncertainty is irreducible (Meyerowitz et al., 1994; Harmanec et al., 1996; Klir & Wierman, 1999; Abellán et al., 2006; Sale et al., 2023; Wimmer et al., 2023). To our knowledge, existing partial-label learning approaches, for example, (Hüllermeier & Beringer, 2005; Liu & Dietterich, 2012; Zhang & Yu, 2015; Ni et al., 2021; Xu et al., 2023), model label information using standard probability, which does not allow to discern aleatoric and epistemic uncertainty.
We propose a partial-label learning approach based on Dempster-Shafer theory that keeps both uncertainties separate and achieves a significantly better trade-off of the number and accuracy of non-rejected predictions when compared to the state-of-the-art.
@@
  • We introduce Dst-Pll, a novel partial-label learning algorithm with a reject option that learns from ambiguously labeled data. The algorithm effectively determines whether to reject a prediction and provides the best trade-off between the number and accuracy of non-rejected predictions.
@@
  • Extensive experiments on artificial and real-world data support our claims. We make our code and data openly available. [^1] https://github.com/anon1248/pll-with-a-reject-option.
@@
  • We analyze Dst-Pll, give a closed-form expression of its expected decision boundary, and prove its risk consistency. We believe these results are of independent interest. The runtime analysis shows that the proposed method is dominated by $k$ nearest-neighbor search, which has an average time complexity of $\mathcal{O}(dk\log n)$ with $d$ features and $n$ training instances.
@@
Structure of the paper. We define the problem setting in Section 2 and Section 3 discusses related work. We propose our method in Section 4, Section 5 features experiments and Section 6 concludes. All proofs and additional experiments are in the appendices.
@@ 2 Problem Statement and Notations
This section formally defines the partial-label learning problem, establishes notations used throughout this work, introduces reject options, and briefly summarizes Dempster-Shafer theory.
@@
Partial-Label Learning (PLL) and Notations. Let $\operatorname{\mathcal{X}}=\mathbb{R}^{d}$ denote a $d$-dimensional real-valued feature space and $\operatorname{\mathcal{Y}}=[l]:=\{1,\dots,l\}$ the finite set of $3\leq l\in\mathbb{N}$ classes. A partial-label learning training dataset $\mathcal{D}=\{(x_{i},s_{i})\mid i\in[n]\}$ of $n$ instances contains feature vectors $x_{i}\in\operatorname{\mathcal{X}}$ and a set of candidate labels $s_{i}\subseteq\operatorname{\mathcal{Y}}$ for each $i\in[n]$. All instances $i$ have an unknown ground-truth label $y_{i}\in\operatorname{\mathcal{Y}}$, and $y_{i}\in s_{i}$. Further, the candidate labels $s_{i}$ can be partitioned into $s_{i}=\{y_{i}\}\cup z_{i}$ with $y_{i}\notin z_{i}$, that is, $z_{i}\subset\mathcal{Y}$ are the false-positive labels.
@@
The sample space is $\operatorname{\Omega}=\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\times 2^{\operatorname{\mathcal{Y}}}$. It is equipped with the smallest Borel $\sigma$-algebra $\operatorname{\mathcal{B}}(\operatorname{\Omega})$, yielding the measurable space $(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$ with respect to which the set of all probability measures $\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$ and a probability measure $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$ is defined. The random variables $X:\operatorname{\Omega}\rightarrow\operatorname{\mathcal{X}}$, $Y:\operatorname{\Omega}\rightarrow\operatorname{\mathcal{Y}}$, and $S:\operatorname{\Omega}\rightarrow 2^{\operatorname{\mathcal{Y}}}$ govern an instance’s features, true label, and candidate labels, respectively. We denote the corresponding marginal measures by associating the corresponding index to $\operatorname{\mathbb{P}}$, for example, $\operatorname{\mathbb{P}}_{X}$ is the distribution of $X$. $\operatorname{\mathbb{P}}_{XY}$ denotes the joint distribution of $(X,Y)$.
@@
PLL aims to train a classifier $g:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$ that, given a loss function $\operatorname{\mathcal{L}}:\operatorname{\mathcal{Y}}\times 2^{\operatorname{\mathcal{Y}}}\to\mathbb{R}_{\geq 0}$, minimizes the risk $\operatorname{\mathcal{R}}(g)=\operatorname{\mathbb{E}}_{(X,S)\sim\operatorname{\mathbb{P}}_{XS}}\operatorname{\mathcal{L}}(g(X),S)$. Common choices for $\operatorname{\mathcal{L}}$ include the average loss (Nguyen & Caruana, 2008; Cour et al., 2011) and the minimum loss (Lv et al., 2020; Xu et al., 2021). The empirical version of the risk substitutes the expectation with a sample mean.
@@
Reject Option. A reject option $\Gamma_{g}:\operatorname{\mathcal{X}}\to 2^{\operatorname{\mathcal{Y}}}$ associated with a trained classifier $g$ either returns $g$’s prediction (accept) or abstains from making any prediction at all (reject), that is, $\Gamma_{g}(x)\in\{\emptyset,\{g(x)\}\}$ for $x\in\operatorname{\mathcal{X}}$ with $\Gamma_{g}(x)=\emptyset$ denoting a reject. We then define the rejection rate $\operatorname{r}(\Gamma_{g})=\operatorname{\mathbb{P}}_{X}(\Gamma_{g}(X)=\emptyset)$ and the expected error $\operatorname{err}(\Gamma_{g})=\operatorname{\mathbb{E}}_{(X,S)\sim\operatorname{\mathbb{P}}_{XS}}[\operatorname{\mathcal{L}}(g(X),S)\mathds{1}_{\{\Gamma_{g}(X)\neq\emptyset\}}]$ of accepted predictions. Naturally, a trade-off arises between the number and accuracy of accepted predictions, which leads us to the risk
@@
for $\lambda\geq 0$. $\operatorname{\mathcal{L}}$ characterizes the cost of misclassification and parameter $\lambda$ the cost of rejecting a prediction. This setting has already been considered in supervised classification (Hellman, 1970; Herbei & Wegkamp, 2006) and in regression (Zaoui et al., 2020). We are the first to study it in PLL.
[…]
$\operatorname{bel}_{\operatorname{m}}(A):=\sum_{B\subseteq A}\operatorname{m}(B)\quad\text{and}\quad\operatorname{pl}_{\operatorname{m}}(A):=\sum_{B\subseteq\operatorname{\mathcal{Y}},A\cap B\neq\emptyset}\operatorname{m}(B)\quad\text{for}\ A\in 2^{\operatorname{\mathcal{Y}}}\text{,}$ (2)
the set of all probability measures supporting $\operatorname{m}$ is
$\operatorname{\mathcal{C}}_{\operatorname{m}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}}):=\big{\{}\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})\mid\operatorname{bel}_{\operatorname{m}}(A)\leq\operatorname{\mathbb{P}}(A)\leq\operatorname{pl}_{\operatorname{m}}(A)\ \text{for all $A\subseteq\operatorname{\mathcal{Y}}$}\big{\}}\text{.}$ (3)
The set $\operatorname{\mathcal{C}}_{\operatorname{m}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})\subseteq\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ is called a credal set (Abellán et al., 2006; Cuzzolin, 2021). Further, DST provides rules to combine $\operatorname{m}$-s from multiple sources (Dempster, 1967; Yager, 1987a, b).
3 Related Work
Imperfect data often renders the application of supervised methods challenging. Weakly supervised learning tackles this setting and encompasses a variety of problem formulations (Bylander, 1994; Hady & Schwenker, 2013; Ishida et al., 2019) including PLL. We discuss related work regarding PLL in Section 3.1. Section 3.2 elaborates on related work in dealing with uncertainty in ML.
3.1 Partial-Label Learning
The early PLL approaches transfer common supervised learning frameworks to the PLL context: Grandvalet (2002) proposes a logistic regression formulation, Jin & Ghahramani (2002) propose an expectation-maximization strategy, Hüllermeier & Beringer (2005) propose a $k$-nearest-neighbors method, Nguyen & Caruana (2008) propose an extension of support-vector machines, and Cour et al. (2011) introduce an average loss formulation allowing for the use of any supervised method. However, these approaches (i) cannot model the relevance of candidate labels in the labeling process or (ii) are not robust to non-uniform noise in the candidate sets. We re-emphasize that (i) and (ii) stem from two different sources of uncertainty: (i) arises from the lack of knowledge about the label relevancies of the PLL model (epistemic uncertainty) and (ii) arises from the random label noise in the candidate sets and is inherent to the PLL problem (aleatoric uncertainty).
More recent approaches address (i) and (ii) in the following ways: Zhang & Yu (2015); Zhang et al. (2016); Xu et al. (2019); Wang et al. (2019); Feng & An (2019); Ni et al. (2021) leverage ideas from representation learning (Bengio et al., 2013), Yu & Zhang (2017); Wang et al. (2019); Feng & An (2019); Ni et al. (2021) extend the maximum-margin idea, Liu & Dietterich (2012); Lv et al. (2020) propose extensions of the expectation-maximization strategy, Zhang et al. (2017); Tang & Zhang (2017); Wu & Zhang (2018) propose stacking and boosting ensembles, Lv et al. (2020); Cabannes et al. (2020) introduce a minimum loss formulation, and Feng et al. (2020); Lv et al. (2020); Xu et al. (2021); Wang et al. (2022); He et al. (2022); Xu et al. (2023); Fan et al. (2024) employ deep-learning techniques such as auto-encoders, data augmentation, and transformers in the context of PLL. These methods address (i) and (ii), featuring already good performance. However, considering a reject option of those methods based on a confidence threshold, as proposed in the supervised setting (Herbei & Wegkamp, 2006), yields sub-optimal results as all of the listed methods mix epistemic and aleatoric uncertainty stemming from (i) and (ii), respectively. We claim that keeping both separate provides a better trade-off for the number and accuracy of non-rejected predictions (Section 5).
3.2 Dealing with Uncertainty

### v4

@@ Abstract
In real-world applications, one often encounters ambiguously labeled data, where different annotators assign conflicting class labels. Partial-label learning allows training classifiers in this weakly supervised setting, where state-of-the-art methods already show good predictive performance. However, even the best algorithms give incorrect predictions, which can have severe consequences when they impact actions or decisions. We propose a novel risk-consistent nearest-neighbor-based partial-label learning algorithm with a reject option, that is, the algorithm can reject unsure predictions. Extensive experiments on artificial and real-world datasets show that our method provides the best trade-off between the number and accuracy of non-rejected predictions when compared to our competitors, which use confidence thresholds for rejecting unsure predictions. When evaluated without the reject option, our nearest-neighbor-based approach also achieves competitive prediction performance.
@@ 1 Introduction
However, even the best algorithms give incorrect predictions. These errors can have severe consequences when they impact actions or decisions. Consider, for example, safety-critical domains such as the classification of medical images (Yang et al., 2009; Lambrou et al., 2011; Senge et al., 2014; Kendall & Gal, 2017; Reamaroon et al., 2019) or the control of self-driving cars (Xu et al., 2014; Varshney & Alemzadeh, 2017; Hubmann et al., 2017; Shafaei et al., 2018; Michelmore et al., 2020). One option to limit fallacies is to employ so-called reject options, which allow one to abstain from certain predictions if unsure and, instead, let humans decide on the label of an instance or the actions to take (Mozannar et al., 2023). Naturally, there is a trade-off arising between the number and accuracy of non-rejected predictions. In the supervised setting, reject options have already been studied, both, for multi-class classification (Charoenphakdee et al., 2021; Cao et al., 2022; Mao et al., 2024; Narasimhan et al., 2024) and regression tasks (Zaoui et al., 2020; Cheng et al., 2023).
In the weakly supervised PLL setting, obtaining sensible reject options is more challenging than in the supervised case as ground truth is not available. Still, a reject option allows for mitigating misclassifications and improving prediction quality. Determining whether to reject a prediction of a PLL algorithm is difficult as the uncertainty involved in prediction-making originates from multiple sources (Hora, 1996; Sale et al., 2023; Wimmer et al., 2023). On the one hand, there is inherent uncertainty due to ambiguously labeled data (aleatoric uncertainty). This ambiguity is influenced by both instances and classes: In some regions of the instance space, annotators are more likely to mislabel instances than in others. Additionally, some class labels are inherently more similar than others. On the other hand, there is uncertainty due to the lack of knowledge regarding the relevance of each candidate label (epistemic uncertainty). To the best of our knowledge, we are the first to study reject options in the context of PLL.
We propose a novel partial-label learning approach based on Dempster-Shafer theory (DST; Dempster 1967; Shafer 1986). In contrast to existing PLL approaches, which use point estimates of the candidate label weights (Hüllermeier & Beringer, 2005; Liu & Dietterich, 2012; Zhang & Yu, 2015; Ni et al., 2021; Xu et al., 2023), our method maintains a feasible region, also known as credal set. Maintaining such a credal set is beneficial in assessing the certainty of predictions as our experiments show.
@@
  • Algorithm with reject option. We introduce Dst-Pll, a novel nearest-neighbor-based partial-label learning algorithm with a reject option that learns from ambiguously labeled data. The algorithm effectively determines whether to reject a prediction and provides the best trade-off between the number and accuracy of non-rejected predictions when compared to the state-of-the-art PLL methods.
@@
  • Experiments. Extensive experiments on artificial and real-world data support our claims. We make our code and data openly available. [^1] https://github.com/mathefuchs/pll-with-a-reject-option.
@@
  • Theoretical analysis. We analyze Dst-Pll, give a closed-form expression of its expected decision boundary under mild assumptions, and prove its risk consistency. The runtime analysis shows that the proposed method’s runtime is dominated by $k$-nearest-neighbor search, which has an average time complexity of $\mathcal{O}(dk\log n)$, with $d$ features and $n$ training instances.
@@
Structure of the paper. We discuss related work in Section 2, define the problem setting in Section 3, and propose our method in Section 4. Section 5 features experiments and Section 6 concludes. All proofs and additional experiments are in the appendices.
2 Related Work
Imperfect data often renders the application of supervised methods challenging. Weakly supervised learning tackles this setting and encompasses a variety of problem formulations (Bylander, 1994; Hady & Schwenker, 2013; Ishida et al., 2019) including PLL. We discuss related work regarding PLL in Section 2.1. Section 2.2 elaborates on related work in dealing with uncertainty in ML through the lens of reject options.
2.1 Partial-Label Learning
The early PLL approaches transfer common supervised learning frameworks to the PLL context: Grandvalet (2002) proposes a logistic regression formulation, Jin & Ghahramani (2002) propose an expectation-maximization strategy, Hüllermeier & Beringer (2005) propose a $k$-nearest-neighbors method, Nguyen & Caruana (2008) propose an extension of support-vector machines, and Cour et al. (2011) introduce an average loss formulation allowing for the use of any supervised method. However, these approaches (i) cannot model the relevance of candidate labels in the labeling process or (ii) are not robust to non-uniform noise in the candidate sets. We emphasize that (i) and (ii) stem from two different sources of uncertainty: (i) arises from the lack of knowledge about the candidate label relevancies of the PLL model (epistemic uncertainty) and (ii) arises from the random label noise in the candidate sets and is inherent to the PLL problem (aleatoric uncertainty).
More recent approaches address (i) and (ii) in the following ways: Zhang & Yu (2015); Zhang et al. (2016); Xu et al. (2019); Wang et al. (2019); Feng & An (2019); Ni et al. (2021) leverage ideas from representation learning (Bengio et al., 2013), Yu & Zhang (2017); Wang et al. (2019); Feng & An (2019); Ni et al. (2021) extend the maximum-margin idea, Liu & Dietterich (2012); Lv et al. (2020) propose extensions of the expectation-maximization strategy, Zhang et al. (2017); Tang & Zhang (2017); Wu & Zhang (2018) propose stacking and boosting ensembles for more robust prediction-making, Lv et al. (2020); Cabannes et al. (2020) introduce a minimum loss formulation to iteratively remove noise, and Feng et al. (2020); Lv et al. (2020); Xu et al. (2021); Wang et al. (2022); He et al. (2022); Xu et al. (2023); Fan et al. (2024) employ deep-learning techniques such as auto-encoders, data augmentation, and transformers in the context of PLL. These methods address (i) and (ii), featuring already good performance. However, considering a reject option of those methods based on a confidence threshold, as proposed in the supervised setting (Ni et al., 2019), yields sub-optimal results as all of the listed methods mix uncertainty stemming from (i) and (ii). We claim that keeping both separate provides a better trade-off for the number and accuracy of non-rejected predictions (Section 5).
In the PLL context, we consider (Hüllermeier & Beringer, 2005; Zhang & Yu, 2015; Zhang et al., 2016; Xu et al., 2019; Wang et al., 2019) as most closely related to our proposed method as these approaches consider an instance’s neighborhood to infer its class label. To the best of our knowledge, we are the first to propose an extension of the $k$-NN classifier that keeps uncertainty from (i) and (ii) separate by leveraging Dempster-Shafer theory.
2.2 Reject Options
Recently, much attention has been given to the study of reject options (Mozannar et al., 2023; Mao et al., 2024; Narasimhan et al., 2024). A reject option allows one to abstain from predictions and defer them to humans rather than making uncertain and possibly harmful decisions.
There are two common strategies in rejecting predictions in the supervised setting: The confidence-based and the classifier-rejector approach (Ni et al., 2019; Cao et al., 2022). The confidence-based strategy uses a threshold on the models’ confidence in order to accept or reject predictions. Common model choices for quantifying the confidences are Bayesian methods (Kingma & Welling, 2014; Kendall & Gal, 2017) and ensembles (Lakshminarayanan et al., 2017; Wimmer et al., 2023). In contrast, the classifier-rejector approach jointly learns the classifier and rejector (Ni et al., 2019; Mao et al., 2024), which can have beneficial theoretical properties. However, the classifier-rejector approach is less flexible than the confidence-based strategy as it is coupled with the concrete loss formulation of the classifier. We propose an extension of the confidence-based strategy for partial-label learning in Section 4.2.
Calibration methods (Naeini et al., 2015; Guo et al., 2017; Ao et al., 2023) are also related to the confidence-based rejection strategy as both are used to make statements about the certainty of predictions. While reject options provide a binary decision, calibration methods modify the predicted confidences such that they align with the observed accuracies. In this sense, both approaches are orthogonal and cannot directly be compared. In our work, we focus on reject options.
@@ 3 Problem Statement and Notations
This section formally defines the partial-label learning problem with reject option, establishes notations used throughout this work, and summarizes Dempster-Shafer theory.
3.1 Partial-Label Learning (PLL) with Reject Option
@@
Let $\operatorname{\mathcal{X}}=\mathbb{R}^{d}$ denote a $d$-dimensional real-valued feature space and $\operatorname{\mathcal{Y}}=[l]:=\{1,\dots,l\}$ the finite set of $3\leq l\in\mathbb{N}$ classes. A partial-label learning training dataset $\mathcal{D}=\{(x_{i},s_{i})\mid i\in[n]\}$ of $n$ instances contains feature vectors $x_{i}\in\operatorname{\mathcal{X}}$ and a set of candidate labels $s_{i}\subseteq\operatorname{\mathcal{Y}}$ for each $i\in[n]$. All instances $i$ have an unknown ground-truth label $y_{i}\in\operatorname{\mathcal{Y}}$, and $y_{i}\in s_{i}$. Further, the candidate labels $s_{i}$ can be partitioned into $s_{i}=\{y_{i}\}\cup z_{i}$ with $y_{i}\notin z_{i}$, that is, $z_{i}\subseteq\mathcal{Y}\setminus\{y_{i}\}$ are the false-positive labels.
@@
The sample space is $\operatorname{\Omega}=\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\times 2^{\operatorname{\mathcal{Y}}}$. It is equipped with the Borel $\sigma$-algebra $\operatorname{\mathcal{B}}(\operatorname{\Omega})$, yielding the measurable space $(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$ with respect to which the set of all probability measures $\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$ is defined. Let $(X,Y,S)\sim\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$, where the random variables $X:\operatorname{\Omega}\rightarrow\operatorname{\mathcal{X}}$, $Y:\operatorname{\Omega}\rightarrow\operatorname{\mathcal{Y}}$, and $S:\operatorname{\Omega}\rightarrow 2^{\operatorname{\mathcal{Y}}}$ govern an instance’s features, true label, and candidate labels, respectively. We denote the corresponding marginal measures by associating the corresponding index to $\operatorname{\mathbb{P}}$, for example, $\operatorname{\mathbb{P}}_{X}$ is the distribution of $X$. $\operatorname{\mathbb{P}}_{XY}$ denotes the joint distribution of $(X,Y)$.
@@
PLL aims to train a classifier $g:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$ that, given a measurable loss function $\operatorname{\mathcal{L}}:\operatorname{\mathcal{Y}}\times 2^{\operatorname{\mathcal{Y}}}\to\mathbb{R}_{\geq 0}$, minimizes the risk $\operatorname{\mathcal{R}}(g)=\operatorname{\mathbb{E}}_{(X,S)\sim\operatorname{\mathbb{P}}_{XS}}\operatorname{\mathcal{L}}(g(X),S)$. Common choices for $\operatorname{\mathcal{L}}$ include the average loss (Nguyen & Caruana, 2008; Cour et al., 2011) and the minimum loss (Lv et al., 2020; Xu et al., 2021). The empirical version of the risk substitutes the expectation with a sample mean, that is, ^$\hat{\operatorname{\mathcal{R}}}(g)=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathcal{L}}(g(x_{i}),s_{i})$.
@@
As misclassifications can be quite harmful, we look at the possibility of rejecting predictions, that is, abstaining from making these predictions and, instead, deferring the decisions to humans. In the PLL setting, a reject option $\Gamma_{g}:\operatorname{\mathcal{X}}\to 2^{\operatorname{\mathcal{Y}}}$ associated with a trained classifier $g$ either returns $g$’s prediction (accept) or abstains from making any prediction at all (reject), that is, $\Gamma_{g}(x)\in\{\emptyset,\{g(x)\}\}$ for $x\in\operatorname{\mathcal{X}}$, with $\Gamma_{g}(x)=\emptyset$ denoting a reject. We then define the rejection probability $\operatorname{r}(\Gamma_{g})=\operatorname{\mathbb{P}}_{X}(\Gamma_{g}(X)=\emptyset)$ and the expected error $\operatorname{err}(\Gamma_{g})=\operatorname{\mathbb{E}}_{(X,S)\sim\operatorname{\mathbb{P}}_{XS}}[\operatorname{\mathcal{L}}(g(X),S)\mathds{1}_{\{\Gamma_{g}(X)\neq\emptyset\}}]$ of accepted predictions. Naturally, a trade-off arises between the number and accuracy of accepted predictions, which leads us to the risk
@@
for $\lambda\geq 0$; $\operatorname{\mathcal{L}}$ characterizes the cost of misclassification and parameter $\lambda$ the cost of rejecting a prediction. In order to make well-informed decisions about when to reject, we are using Dempster-Shafer theory, which we elaborate on in the following.
3.2 Dempster-Shafer Theory (DST)
Dempster-Shafer theory (DST; Dempster 1967; Shafer 1986) allows for dealing with uncertainty by assigning probability mass to sets of events without specifying the probabilities of individual labels; incorrect labels do not obtain any probability mass. DST builds upon two core quantities, so-called belief and plausibility. Informally, belief collects all evidence that supports a hypothesis and plausibility collects all evidence that does not contradict a hypothesis. We argue that DST is a perfect fit for partial-label learning as one may interpret the ambiguous candidate sets as evidence regarding a label hypothesis. We further exploit belief and plausibility to inform our reject option, taking into account the difference between the supporting and non-conflicting evidence. In contrast, existing PLL approaches (Hüllermeier & Beringer, 2005; Cour et al., 2011; Liu & Dietterich, 2012; Zhang & Yu, 2015; Ni et al., 2021; Xu et al., 2023) initially assign some probability mass to each label candidate and subsequently refine them. By doing so, most probability mass is first allocated to labels that are certainly incorrect, as only one candidate is the true label. This known method renders handling the noise coming from incorrect candidate labels challenging.
With these intuitions, we now recall DST formally. In DST, a basic probability assignment (bpa) $\operatorname{m}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1]$ assigns probability mass to subsets of $\operatorname{\mathcal{Y}}$. $\operatorname{m}$ satisfies $\operatorname{m}(\emptyset)=0$ and $\sum_{A\subseteq\operatorname{\mathcal{Y}}}\operatorname{m}(A)=1$. This differs from standard probability as $\operatorname{\mathbb{P}}(\operatorname{\mathcal{Y}})=1$ for any $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ but $\operatorname{m}(\operatorname{\mathcal{Y}})\leq 1$. Also, the mass allocated to non-intersecting sets does not necessarily add up to the mass allocated to the union, that is, one can have $\operatorname{m}(\{1,2\})\neq\operatorname{m}(\{1\})+\operatorname{m}(\{2\})$. In this sense, DST allows for more flexibility as one can allocate mass on the set $\{1,2\}$ without needing to specify any mass for $\{1\}$ and $\{2\}$ if uncertain. The sets $A\subseteq\operatorname{\mathcal{Y}}$ with $\operatorname{m}(A)>0$ are called focal sets of $\operatorname{m}$. The mass allocated to the set of all possible alternatives $\operatorname{m}(\operatorname{\mathcal{Y}})$ can be interpreted as the degree of ignorance; it is the mass not supporting a specific alternative within $\operatorname{\mathcal{Y}}$.
The basic probability assignment $\operatorname{m}$ does not induce a single probability measure $\operatorname{\mathbb{P}}$ on $(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ but rather a set of probability measures $\operatorname{\mathcal{C}}_{\operatorname{m}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$, which is called credal set (Abellán et al., 2006; Cuzzolin, 2021). The probability measures $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{m}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ are restricted by imposing lower and upper bounds, which are called belief and plausibility, respectively. They are defined as follows:
$\operatorname{bel}_{\operatorname{m}}(A):=\sum_{B\subseteq A}\operatorname{m}(B)\quad\text{and}\quad\operatorname{pl}_{\operatorname{m}}(A):=1-\operatorname{bel}_{\operatorname{m}}(\operatorname{\mathcal{Y}}\setminus A)=\sum_{B\subseteq\operatorname{\mathcal{Y}},A\cap B\neq\emptyset}\operatorname{m}(B)\quad\text{for}\ A\in 2^{\operatorname{\mathcal{Y}}}\text{.}$ (2)
Recall that belief collects all evidence that supports a hypothesis $A\in 2^{\operatorname{\mathcal{Y}}}$ (or a more specific one $B\subseteq A$), and plausibility collects all evidence that does not contradict a hypothesis $A\in 2^{\operatorname{\mathcal{Y}}}$ (or a more specific one $B\subseteq A$). With belief and plausibility as above, the set of all probability measures supporting $\operatorname{m}$ is
$\operatorname{\mathcal{C}}_{\operatorname{m}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}}):=\big{\{}\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})\mid\operatorname{bel}_{\operatorname{m}}(A)\leq\operatorname{\mathbb{P}}(A)\leq\operatorname{pl}_{\operatorname{m}}(A)\ \text{for all $A\subseteq\operatorname{\mathcal{Y}}$}\big{\}}\subseteq\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})\text{.}$ (3)
Further, DST provides rules to combine $\operatorname{m}$-s from multiple sources (Dempster, 1967; Yager, 1987a; b). This is beneficial in the PLL setting as there is several conflicting evidence about the class labels within a neighborhood of instances. Estimating a credal set $\operatorname{\mathcal{C}}_{\operatorname{m}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ from such a neighborhood allows us to construct an effective reject option, which we detail in Section 4.2.